OpenStudy (kimberly101):

Shannon buys a used car for $7500. The dealer requires a 15% deposit. Calculate the deposit Shannon has to pay.

OpenStudy (anonymous):

15/100 *7500=15 *75=the deposit Shannon has to pay=1125

OpenStudy (anonymous):

Kimberly101 understood

OpenStudy (kimberly101):

yes thankyou very much